QUESTIONS Two los, A and B, in Drosophila minerals have two ancies oach UAE and Ray A population shows the toioeng hagiotape recanoe IAL I o2 FD If the frequency of the B allele is 06, that is the frequency of the aihapdepe? OB.01 O CO2 03 OLO .5 o a 2. on Oth 0.7 0.8 OI 0.9 QUESTION 6 What would be the requency of the Athl haplotype in a population won the same alolo tracancies but in linkage equlinum? O Ba1 OC ac2 Oda ce O - 0 0 006 0 0 08 OLOR 0 0 1
Answer #1

Ans) According to the "Wahlund effect",

If the frequency of the B allele is 0.6.

Then the frequency b is = 1 - 0.6 = 0.4

The frequency of A allele is = AB/B = 0.2/0.6 = 0.33

the frequency of "a" allele is 1 - 0.33 = 0.67

Now the frequency of "Ab" allele is 0.33 X 0.4 = 0.133

The frequency of "aB" allele is 0.67 X 0.6 = 0.402

The second question is not clearly readable.